OpenStudy (anonymous):

The governor of state A earns $48,110more than the governor of state B. if the total of their salaries is $304,460. Find the salaries of each

OpenStudy (mathstudent55):

Let one salary be x. The other salary is y. The difference of the salaries is x - y = 48110 The sum of the salaries is x + y = 304460 Solve the system of equations for x and y. x - y = 48110 x + y = 304460

OpenStudy (anonymous):

or... A = B + 48110 A + B = 304460 now sub B + 48110 in for A in the 2nd equation and solve for B. Once you have B solved, sub that back into either of the equations and solve for A.

OpenStudy (anonymous):

A = 176,285 B = 128,175 Thanks!

OpenStudy (anonymous):

glad to help :)